an inefficient secretary places  5 different letters into 5 different addressed envolopes at random find the probability that atleast one of the letters will arrive proper destination

Grade:12

1 Answers

AskiitianExpert Shine
14 years ago

Hi

the probability that atleast one of the letter goes in the right envelope = 1- (probability that none of the letters go in the right envelope)
now, total ways = 5! = 120
ways that none of the letters go in the right envelope = 4! = 24
probability that none of the letters go in the right envelope = 24/120 = 1/5
probability that atleast one of the letter goes in the right envelope = 1 - 1/5 = 4/5